# Spectral Convolutional Conditional Neural Processes

### v4

@@ 3 Spectral Convolutional Conditional Neural Process
This locality constraint limits the model’s capacity to capture long-range dependencies and to integrate information from observations distributed across large domains (Peng_2017_CVPR; wang2018non; ramachandran2019stand; wang2020axial). The issue becomes particularly pronounced when handling sparse or irregularly sampled data, where effective global reasoning cannot emerge solely from local convolution operations. While increasing the kernel size could enlarge the receptive field, it leads to a rapid growth in both parameter count and computational cost. Transformer-based architectures offer an alternative by enabling global interactions (vaswani2017attention); however, they typically incur quadratic rather than linear complexity in the number of inputs (nguyen2022transformer; feng2022latent; ashman2024translation; ashman2025gridded).
@@
To overcome this limitation without relying on prohibitively large filters or expensive transformers, we exploit the Fourier representation of signals. This is motivated by the well-established observation that many natural signals are approximately band-limited (see Section 2.1), implying that their Fourier representation offers a more compact encoding relative to its physical-domain counterpart. Concretely, we instantiate $\varphi_{d}$ via spectral convolution modules based on equation 1. This substitution effectively enlarges the receptive field, enabling the model to capture global structures from sparse or irregularly sampled data without incurring a parameter count explosion. We refer to the resulting model as the Spectral Convolutional Conditional Neural Process (SConvCNP).
@@
The computational cost of the SConvCNP comprises three parts: (i) $\mathcal{O}(|\mathcal{D}_{c}||\mathcal{G}|)$ to compute the discretized functional embedding on the grid $\mathcal{G}$ (equation 3); (ii) $\mathcal{O}(|\mathcal{G}|\log|\mathcal{G}|)$ for the FFT-based spectral convolution (equation 1); and (iii) $\mathcal{O}(|\mathcal{D}_{q}||\mathcal{G}|)$ to interpolate grid embeddings at the query locations (equation 4). Overall, the total complexity is $\mathcal{O}\bigl(|\mathcal{G}|(|\mathcal{D}_{c}|+\log|\mathcal{G}|+|\mathcal{D}_{q}|)\bigr)$, matching the ConvCNP complexity of $\mathcal{O}\bigl(|\mathcal{G}|(|\mathcal{D}_{c}|+1+|\mathcal{D}_{q}|)\bigr)$ up to logarithmic factors. Both architectures therefore scale linearly with task size. A key limitation, however, is that $|\mathcal{G}|$ grows exponentially with input dimension, restricting these methods to low-dimensional domains. In contrast, transformer-based NPs (TNPs; kim2019attentive; nguyen2022transformer; feng2022latent; ashman2024translation) avoid gridding entirely and thus scale more gracefully with input dimension. Instead, they scale quadraticaly in task size $\mathcal{O}(|\mathcal{D}_{c}|^{2}+|\mathcal{D}_{c}||\mathcal{D}_{q}|)$.
@@
Although this addition breaks translation equivariance, it helps improving the performance (see Section B.5 for ablation results). Future work may explore relative positional encodings (shaw2018self; su2024roformer), which preserve translation equivariance while providing spatial context.
@@
Unlike the continuous Fourier transform, the DFT—and therefore the FFT—depends inherently on the grid $\mathcal{G}$ on which $\varphi_{e}[\mathcal{D}_{c}]$ (or $\widetilde{\varphi}_{e}[\mathcal{D}_{c}]$) is represented. This dependence arises from both the grid resolution and its physical extent: changing either alters the resulting Fourier coefficients and can lead to inconsistent behavior in the outputs (see Section A). Sensitivity to resolution is not unique to DFT; CNNs exhibit analogous issues (raonic2023convolutional; bartolucci2023representation). For example, ConvCNP mitigates this effect by fixing the grid resolution. While spatial CNNs become stable once the resolution is fixed, DFT-based methods remain sensitive unless both the resolution and the physical range are controlled. Accordingly, we fix both, choosing a domain sufficiently large to cover all context and query inputs across tasks. When this is impractical, the domain can instead be divided into (possibly overlapping) fixed-size patches, with spectral convolutions applied independently to each patch and the outputs aggregated—mirroring the mechanism of standard convolution layers. Related ideas in efficient transformer architectures suggest that this is a promising direction for future work (beltagy2020longformer; zaheer2020big; liu2021swin; ding2023longnet).
@@ 4 Experiments
Figure 1: Example predictions for synthetic data. For each model, the blue curve shows its predictive mean and the shaded region indicates $\pm 2$ standard deviations under the corresponding Gaussian predictive distribution. In the first two rows, where the data are generated from a Gaussian process, the true joint density is shown in purple: the dash-dotted line marks the true mean, and the shaded band spans $\pm 2$ standard deviations around that mean. Red points denote query samples drawn from the ground-truth process, and black points indicate the context observations.
We evaluate our framework on four regression benchmarks and compare its performance against several members of CNPs family. Specifically, we include the original CNP (garnelo2018conditional), the Attentive CNP (AttCNP; kim2019attentive), the Convolutional CNP (ConvCNP; gordon2019convolutional), the Transformer Neural Process (TNP; nguyen2022transformer), and the Translation-Equivariant Transformer Neural Process (TETNP; ashman2024translation). To evaluate model performance, we report both the log-likelihood and root-mean-squared error (RMSE) metrics. Additional details on datasets, architectures, and training procedures are provided in Appendix B. The implementation and experimental code are available at https://github.com/peiman-m/SConvCNP.
@@ 4.1 Synthetic 1-D Regression
Table 1: Comparison of predictive performance across methods on synthetically generated tasks (4 seeds). Lower RMSE and higher log-likelihood indicate better performance. For each metric and experimental setting, boldface highlights the models ranking in the top two.
@@
We begin by evaluating models on four synthetic benchmarks generated from distinct stochastic processes: a GP with a Matérn–5/2 kernel, a GP with a periodic kernel, a sawtooth-wave generator, and a square-wave generator. For each benchmark, the parameters of the generative process—kernel hyperparameters for the GPs, frequency and direction for the sawtooth generator, and frequency and duty cycle for the square-wave generator—are sampled randomly. Complete experimental details are provided in Appendix B.1. Table 1 reports the average evaluation metrics over 1,000 test batches, each containing 16 tasks (each task defined in equation 2). For every task—independent of batch—both $x_{c}$ and $x_{q}$ are sampled independently from $\mathcal{U}[-3,3)$. For each test batch, the number of context points, shared by all tasks within that batch, is drawn independently according to $|\mathcal{I}_{c}|\sim\mathcal{U}[5,25)$. The number of query points, however, is fixed at $|\mathcal{I}_{q}|=256$ for all the test tasks.
As shown, SConvCNP consistently outperforms or closely matches strong baselines, including TETNP and ConvCNP, which represent the current state of the art. Figure 1 provides qualitative comparisons of predictive maps produced by SConvCNP, ConvCNP, and TETNP (other models are omitted due to space constraints).
@@ 4.2 Predator–Prey Dynamics
$dU_{t}=\alpha U_{t}\,dt-\beta U_{t}V_{t}\,dt+\sigma U_{t}^{\nu}\,dB^{(1)}_{t},\qquad dV_{t}=-\gamma V_{t}\,dt+\delta U_{t}V_{t}\,dt+\sigma V_{t}^{\nu}\,dB^{(2)}_{t}$,
where $B^{(1)}_{t}$ and $B^{(2)}_{t}$ are independent Brownian motions. In the deterministic part of the model, the prey population grows exponentially at rate $\alpha$ in the absence of predation, while the predator population declines at rate $\gamma$ without access to prey. The interaction terms $\beta U_{t}V_{t}$ and $\delta U_{t}V_{t}$ describe predation and the resulting reproductive gain for predators. Stochasticity is introduced through the multiplicative noise terms $\sigma U_{t}^{\nu}$ and $\sigma V_{t}^{\nu}$, where $\sigma$ controls the overall noise intensity and $\nu$ determines how fluctuations scale with population size.
@@
Table 2: Comparison of predictive performance across methods on test tasks constructed from the Pred–Prey population dynamics simulation (4 seeds). Lower RMSE and higher log-likelihood indicate better performance. Boldface highlights the models ranking in the top two for each metric.
@@
Figure 2: Illustrative model predictions based on simulated trajectories from Lotka–Volterra predator–prey dynamics. Black points represent the context observations. For each model, the blue curve depicts the predictive mean, and the shaded region marks the $\pm 2$-standard-deviation band of the associated Gaussian predictive distribution. Ground-truth traffic measurements, shown in red, are included as a reference.
@@ 4.3 Traffic Flow
Table 3: Comparison of predictive performance across methods on test tasks constructed from California traffic flow measurements (4 seeds). Lower RMSE and higher log-likelihood indicate better performance. Boldface highlights the models ranking in the top two for each metric.
@@
Figure 3: Illustrative model predictions on the California traffic-flow dataset. Black points represent the context observations. For each model, the blue curve depicts the predictive mean, and the shaded region marks the $\pm 2$-standard-deviation band of the associated Gaussian predictive distribution. Ground-truth traffic measurements, shown in red, are included as a reference.
@@ 4.4 Image Completion
For our final experiment, we evaluate model performance on an image-completion task formulated as a spatial regression problem, where the model maps 2D pixel coordinates to their corresponding intensity values. We use images from the Describable Textures Dataset (DTD; cimpoi2014describing), and construct each task from a processed $64\times 64$ subsampled crop of an original image. Because each task contains a large number of context and query pixels, we were unable to fit TETNP within our computational budget—even after reducing its size—so we exclude it from this experiment. Additional experimental details are provided in Appendix B.4. Table 4 reports results averaged over 1,880 test tasks, evaluated in batches of 16. For each batch, the number of context points—shared across all tasks—is sampled as $|\mathcal{I}_{c}|\sim\mathcal{U}[5,1024)$; all remaining pixels serve as query points, so that $|\mathcal{I}_{q}|=64\times 64-|\mathcal{I}_{c}|$. Among the baselines, SConvCNP achieves the highest log-likelihoods and the second-lowest RMSE. Figure 4 presents qualitative comparisons of the predictive maps produced by the top three models.
@@
Figure 4: Illustrative model outputs for image-completion tasks using samples from the DTD dataset. Grey pixels indicate query regions, and the remaining pixels serve as context observations. For each model, query pixels are filled with the mean of the predictive distribution.
Table 4: Comparison of predictive performance across methods on image-completion test tasks constructed from images of the DTD dataset (4 seeds). Lower RMSE and higher log-likelihood indicate better performance. Boldface highlights the models ranking in the top two for each metric.
@@ 5.1 Neural PDE Solvers
5.2 Function Space Inference
In non-parametric Bayesian modeling, GPs and deep GPs (damianou2013deep) provide flexible function-space priors with well-calibrated uncertainty estimates, but their computational cost becomes prohibitive on large datasets. This challenge has motivated the development of Bayesian neural networks (BNNs; mackay1992practical; hinton1993keeping; neal2012bayesian), which combine neural network expressiveness with Bayesian uncertainty quantification. However, specifying meaningful priors over network weights remains notoriously difficult. Recent work therefore reframes Bayesian inference in neural networks as learning a posterior over the functions induced by stochastic weights (wolpert1993bayesian; qiu2023should). Variational implicit processes (VIPs; ma2019variational; santana2021function; ortega2022deep) generalize GPs by defining implicit stochastic processes through latent variables, while functional variational BNNs (fBNNs; sun2019functional) enforce alignment between BNN-induced priors and target priors by minimizing a functional KL divergence—though this objective can be difficult to compute and, in some cases, ill-posed (burt2020understanding). Subsequent work has aimed to address these limitations (ma2021functional; rudner2022tractable; wild2022generalized; rudner2023function; wu2024functional). More recently, several approaches have extended diffusion models to function spaces (phillips2022spectral; dutordoir2023neural; mathieu2023geometric; franzese2023continuous; lim2023score).
@@ 6 Conclusion
In this work, we introduced the Spectral Convolutional Conditional Neural Process (SConvCNP), a new member of the CNP family that leverages advances in operator learning—particularly the Fourier Neural Operator—to model probabilistic prediction maps within the neural processes framework, with a special emphasis on the Convolutional Conditional Neural Process (ConvCNP). Through experiments on both synthetic and real datasets, we show that SConvCNP delivers highly competitive performance relative to recent state-of-the-art methods, highlighting the promise of bridging research on neural processes with developments in operator learning.

### v5

@@ 3 Spectral Convolutional Conditional Neural Processes
The decoder $\varphi_{d}$ in ConvCNPs is typically parameterized using standard CNNs such as U-Net (ronneberger2015u) or ResNet (he2016deep). These architectures employ discrete convolutional kernels—finite sets of learnable parameters that define localized filters operating over neighboring grid points. The kernel size, fixed a priori, determines the receptive field of each convolution (ding2022scaling) and is generally much smaller than the overall extent of the input signals in the physical domain (romero2021ckconv; knigge2023modelling). This locality constraint fundamentally limits a model’s ability to capture long-range dependencies and to integrate information from observations distributed across large spatial or temporal domains (Peng_2017_CVPR; wang2018non; ramachandran2019stand; wang2020axial). The issue becomes particularly pronounced when handling sparse or irregularly sampled data, where effective global reasoning cannot emerge solely from local convolution operations. Although enlarging the convolutional kernel increases the receptive field, it results in a rapid escalation of both parameter count and computational cost. Transformer-based architectures mitigate this issue by enabling explicit global interactions (vaswani2017attention); however, they incur quadratic, rather than linear, complexity in the number of inputs, rendering them impractical for large context sets unless one resorts to approximation schemes (nguyen2022transformer; feng2022latent; ashman2024translation; ashman2025gridded).
@@
To overcome this limitation without relying on prohibitively large filters or expensive transformers, we exploit the Fourier representation of signals. This is motivated by the well-established observation that many natural signals are approximately band-limited (see Section 2.1), implying that their Fourier representation offers a more compact encoding relative to its physical-domain counterpart. Concretely, we instantiate $\varphi_{d}$ via spectral convolution modules based on equation 1. This substitution effectively enlarges the receptive field, enabling the model to capture global structures from sparse or irregularly sampled data without incurring a parameter count explosion. We refer to the resulting models as spectral convolutional conditional neural processes (SConvCNPs).
@@
The computational cost of SConvCNPs comprises three parts: (i) $\mathcal{O}(|\mathcal{D}_{c}||\mathcal{G}|)$ to compute the discretized functional embedding on the grid $\mathcal{G}$ (equation 3); (ii) $\mathcal{O}(|\mathcal{G}|\log|\mathcal{G}|)$ for the FFT-based spectral convolutions (equation 1); and (iii) $\mathcal{O}(|\mathcal{D}_{q}||\mathcal{G}|)$ to interpolate grid embeddings at the query locations (equation 4). Overall, the total complexity is $\mathcal{O}\bigl(|\mathcal{G}|(|\mathcal{D}_{c}|+\log|\mathcal{G}|+|\mathcal{D}_{q}|)\bigr)$, matching the $\mathcal{O}\bigl(|\mathcal{G}|(|\mathcal{D}_{c}|+1+|\mathcal{D}_{q}|)\bigr)$ complexity of ConvCNPs up to logarithmic factors. Both architectures therefore scale linearly with task size. A key limitation, however, is that $|\mathcal{G}|$ grows exponentially with the input dimension $d_{x}$, restricting these methods to low-dimensional domains. In contrast, transformer-based NPs (TNPs; kim2019attentive; nguyen2022transformer; feng2022latent; ashman2024translation) avoid gridding entirely and thus scale more gracefully with input dimensionality. Instead, their computational cost scales quadratically with the task size, typically as $\mathcal{O}(|\mathcal{D}_{c}|^{2}+|\mathcal{D}_{q}|^{2})$, with minor variations depending on the specific implementation (see Section B.1.1).
@@
While this augmentation explicitly breaks translation equivariance, we empirically observe performance improvements (see Section B.5). An interesting direction for future work is to investigate relative positional encodings (shaw2018self; su2024roformer), which can provide spatial context while preserving translation equivariance.
@@
Unlike the continuous Fourier transform, the DFT—and therefore the FFT—depends inherently on the grid $\mathcal{G}$ on which $\varphi_{e}[\mathcal{D}_{c}]$ (or $\widetilde{\varphi}_{e}[\mathcal{D}_{c}]$) is represented. This dependence arises from both the grid resolution and its physical extent: changing either alters the resulting Fourier coefficients and can lead to inconsistent behavior in the outputs (see Section A). Sensitivity to resolution is not unique to DFT; CNNs exhibit analogous issues (raonic2023convolutional; bartolucci2023representation). For example, ConvCNPs mitigate this effect by fixing the grid resolution. While spatial CNNs become stable once the resolution is fixed, DFT-based methods remain sensitive unless both the resolution and the physical range are controlled. Accordingly, we fix both, choosing a domain sufficiently large to cover all context and query inputs across tasks. When this is impractical, the domain can instead be divided into (possibly overlapping) fixed-size patches, with spectral convolutions applied independently to each patch and the outputs aggregated—mirroring the mechanism of standard convolution layers. Related ideas in efficient transformer architectures suggest that this is a promising direction for future work (beltagy2020longformer; zaheer2020big; liu2021swin; ding2023longnet).
@@ 4 Experiments
Figure 1: Example predictions on synthetic data. For each model, the blue curve denotes the predictive mean, and the shaded region corresponds to $\pm 2$ standard deviations of the model’s Gaussian predictive distribution. In the first two rows, where the data are generated from Gaussian processes, the ground-truth distribution is shown in purple: the dash–dotted curve indicates the true mean, and the shaded band represents $\pm 2$ standard deviations around it. Black points denote context observations, and the red point indicates the queries.
We evaluate our framework on four regression benchmarks and compare its performance against several members of CNPs family. Specifically, we include the original CNP (garnelo2018conditional), the Attentive CNP (AttCNP; kim2019attentive), the Convolutional CNP (ConvCNP; gordon2019convolutional), the diagonal variant of the Transformer Neural Process (TNP; nguyen2022transformer), and the Translation-Equivariant Transformer Neural Process (TETNP; ashman2024translation). For each experimental setting, all models are trained using four random seeds. Performance is evaluated using the final aggregated log-likelihood and root-mean-squared error (RMSE), and we report the mean $\pm$ standard deviation across runs. Additional details regarding datasets, architectures, and training procedures are provided in Appendix B. Our implementation and experimental code are publicly available at https://github.com/peiman-m/SConvCNP.
@@ 4.1 Synthetic 1-D Regression
Table 1: Comparison of predictive performance across methods on synthetically generated tasks. Lower RMSE and higher log-likelihood indicate better performance. For each metric and experimental setting, boldface denotes the top-two performing models.
@@
We begin by evaluating models on four synthetic benchmarks generated from distinct stochastic processes: a GP with a Matérn–5/2 kernel, a GP with a periodic kernel, a sawtooth-wave generator, and a square-wave generator. For each benchmark, the parameters of the generative process—kernel hyperparameters for the GPs, frequency and direction for the sawtooth generator, and frequency and duty cycle for the square-wave generator—are sampled randomly. Complete experimental details are provided in Appendix B.1. Table 1 reports the average evaluation metrics over 1,000 test batches, each containing 16 tasks (each task defined in equation 2). For every task—independent of batch—both $x_{c}$ and $x_{q}$ are sampled independently from $\mathcal{U}[-3,3)$. For each test batch, the number of context points, shared by all tasks within that batch, is drawn independently according to $|\mathcal{I}_{c}|\sim\mathcal{U}[5,25)$. The number of query points, however, is fixed at $|\mathcal{I}_{q}|=256$ for all the test tasks. As shown, SConvCNP consistently outperforms or closely matches strong baselines, including TETNP and ConvCNP, which represent the current state of the art. Figure 1 provides qualitative comparisons of predictive maps produced by SConvCNP, ConvCNP, and TETNP.
@@ 4.2 Predator–Prey Dynamics
$dU_{t}=\alpha U_{t}\,dt-\beta U_{t}V_{t}\,dt+\sigma U^{\nu}_{t}\,dB^{(1)}_{t},\qquad dV_{t}=-\gamma U_{t}\,dt+\delta U_{t}V_{t}\,dt+\sigma V^{\nu}_{t}\,dB^{(2)}_{t}$,
where $B^{(1)}_{t}$ and $B^{(2)}_{t}$ are independent Brownian motions.In the deterministic component of the dynamics, $U_{t}$ grows exponentially at rate $\alpha$, while $V_{t}$ decays at rate $\gamma$. The bilinear interaction terms $\beta U_{t}V_{t}$ and $\delta U_{t}V_{t}$ model predation and the corresponding transfer of biomass from prey to predators. To account for stochastic fluctuations commonly observed in empirical population counts, the dynamics are augmented with multiplicative noise terms $\sigma U^{\nu}_{t}dB^{(1)}_{t}$ and $\sigma V^{\nu}_{t}dB^{(2)}_{t}$. Here, $\sigma$ controls the noise magnitude, while $\nu$ determines how the variability scales with population size.
@@
Table 2: Comparison of predictive performance across methods on tasks constructed from the Lotka–Volterra system simulation. Lower RMSE and higher log-likelihood indicate better performance. For each metric and experimental setting, boldface denotes the top-two performing models.
@@
Figure 2: Illustrative predictions on simulated Lotka–Volterra predator–prey trajectories. Black markers denote the context observations. The blue curve shows the predictive mean, while the shaded region corresponds to the $\pm 2$ standard deviation interval of the Gaussian predictive distribution. Ground-truth population trajectories from the simulator are plotted in red.
@@ 4.3 Traffic Flow
Table 3: Comparison of predictive performance across methods on tasks constructed from California traffic flow measurements. Lower RMSE and higher log-likelihood indicate better performance. For each metric and experimental setting, boldface denotes the top-two performing models.
@@
Figure 3: Illustrative predictions on the California traffic-flow dataset. Black markers denote the context observations. The blue curve shows the predictive mean, while the shaded region corresponds to the $\pm 2$ standard deviation interval of the Gaussian predictive distribution. Ground-truth traffic measurements are shown in red.
@@ 4.4 Image Completion
For our final experiment, we evaluate model performance on an image-completion task formulated as a spatial regression problem, where the model maps 2D pixel coordinates to their corresponding intensity values. We use images from the Describable Textures Dataset (DTD; cimpoi2014describing), and construct each task from a processed $64\times 64$ subsampled crop of an original image. Because each task contains a large number of context and query pixels, we were unable to fit TETNP within our computational budget—even after reducing its size—so we exclude it from this experiment. Additional experimental details are provided in Appendix B.4. Table 4 reports results averaged over 1,880 test tasks, evaluated in batches of 16. For each batch, the number of context points—shared across all tasks—is sampled as $|\mathcal{I}_{c}|\sim\mathcal{U}[5,1024)$; all remaining pixels serve as query points, so that $|\mathcal{I}_{q}|=64\times 64-|\mathcal{I}_{c}|$. Among the baselines, SConvCNP achieves the highest log-likelihoods and the second-lowest RMSE. Figure 4 presents qualitative comparisons of the predictive means produced by the top three models.
@@
Figure 4: Illustrative model outputs for image completion on the DTD dataset. Gray pixels indicate query regions, while the remaining pixels serve as context observations. For each model, query regions are completed using the mean of the predictive distribution.
Table 4: Comparison of predictive performance across methods on image-completion tasks constructed from the DTD dataset. Lower RMSE and higher log-likelihood indicate better performance. For each metric and experimental setting, boldface denotes the top-two performing models..
@@ 5 Related Works
Function Space Inference
In non-parametric Bayesian modeling, GPs and deep GPs (damianou2013deep) provide flexible function-space priors with well-calibrated uncertainty estimates, but their computational cost becomes prohibitive on large datasets. This challenge has motivated the development of Bayesian neural networks (BNNs; mackay1992practical; hinton1993keeping; neal2012bayesian), which combine neural network expressiveness with Bayesian uncertainty quantification. However, specifying meaningful priors over network weights remains notoriously difficult. Recent work therefore reframes Bayesian inference in neural networks as learning a posterior over the functions induced by stochastic weights (wolpert1993bayesian; qiu2023should). Variational implicit processes (VIPs; ma2019variational; santana2021function; ortega2022deep) generalize GPs by defining implicit stochastic processes through latent variables, while functional variational BNNs (fBNNs; sun2019functional) enforce alignment between BNN-induced priors and target priors by minimizing a functional KL divergence—though this objective can be difficult to compute and, in some cases, ill-posed (burt2020understanding). Subsequent work has aimed to address these limitations (ma2021functional; rudner2022tractable; wild2022generalized; rudner2023function; wu2024functional). Orthogonally, energy-based models have been explored for representing stochastic processes in function space (yang2020energy; lim2022energy). More recently, several approaches have extended diffusion and flow models to function spaces (phillips2022spectral; dutordoir2023neural; mathieu2023geometric; franzese2023continuous; lim2023score; pmlr-v238-kerrigan24a; zhang2025flow).
@@ 6 Conclusion
In this work, we introduced spectral convolutional conditional neural processes (SConvCNPs), a class of conditional neural processes that incorporates ideas from operator learning—particularly Fourier neural operators—into the neural processes framework, with an emphasis on convolutional conditional neural processes. Empirical evaluations on a collection of synthetic and real-world datasets indicate that SConvCNPs perform comparably to strong baselines. These results suggest that integrating techniques from operator learning into neural processes is a viable direction for probabilistic function modeling and warrants further investigation.